OpenStudy (anonymous):

the capacity of a small childrens swimming pool is 106 gallons of water. there are currently 15 gallons of water in the pool. you are filling the pool with water at a rate of 2 gallons per minute. a. write an equation that give the amount y(in gallon) of water in the pool as a function of the number x of minutes from now. b. after how many minutes will the pool be fill. show work.

OpenStudy (anonymous):

someone help pleaseeeee

OpenStudy (anonymous):

helloo?

OpenStudy (anonymous):

helloooooooooooooooooooooo

OpenStudy (anonymous):

y = 15 +2x

OpenStudy (anonymous):

y= 2x +15 so then 106=2x +15 91=2x x= 45.5 minutes

OpenStudy (anonymous):

is this for a?

OpenStudy (anonymous):

y= 2x+15 is part a

OpenStudy (anonymous):

the rest is b?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank u:)